OpenStudy (anonymous):

Help finding f^-1(x)

OpenStudy (anonymous):

\[f(x)\sqrt[3]{x-8}\]

hartnn (hartnn):

write it like y= \(\sqrt[3]{x-8}\) then cube both sides, what u get ?

OpenStudy (anonymous):

cube each side?

OpenStudy (anonymous):

so f^-1 would look like y = cubed root of x - 8 x = cubed root of y -8 so you would cube both sides x^3 = y-8 add 7 to both sides y = x^3 + 8 ??

hartnn (hartnn):

that seems correct to me. good work :)

OpenStudy (anonymous):

oh wow lol thanks!

hartnn (hartnn):

welcome ^_^